OpenStudy (anonymous):

Find the volume of a sphere with a diameter of 13 inches. Use 3.14 to approximate π. Round answers to the nearest hundredth.

OpenStudy (jdoe0001):

same as before

OpenStudy (anonymous):

O OK

OpenStudy (anonymous):

IM SORRY THE FORMULA DIDNT POP UP WATS THE FORMULA AGAIN

OpenStudy (anonymous):

I am up with 2122.64

OpenStudy (anonymous):

came*

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

keep in mind that the DIAMETER is 13 r = radius

OpenStudy (ranga):

The question seeks Volume of the sphere not the surface area.

OpenStudy (jdoe0001):

ohh shoot... ohh ... I see.. you need VOLUME previouly was surface... ok then volume of a sphere = \(\large \boldsymbol{\cfrac{4}{3}\pi r^3}\)

OpenStudy (anonymous):

formula isnt showing, sorry

OpenStudy (jdoe0001):

hmmm I assume you're using ... Internet Explorer?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

hmm ahemm try in another browser, same page :|

OpenStudy (ranga):

Volume of a sphere = 4/3(pi)r^3

OpenStudy (ranga):

What are you getting for the volume?

OpenStudy (anonymous):

i got 27594.32 :\

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i did it wrong

OpenStudy (ranga):

radius r = diameter / 2 = 13 / 2 = 6.5 V = 4/3(pi)r^3 = 4/3 x 3.14 x 6.5 x 6.5 x 6.5 = ?

OpenStudy (jdoe0001):

\(\bf \cfrac{4}{3}\pi r^3\\ \quad \\ \textit{diameter is 13, thus radius is half that}\\ \quad \\ \cfrac{4}{3}\pi \left(\frac{13}{2}\right)^3\)

OpenStudy (anonymous):

its 1150.346

OpenStudy (jdoe0001):

yeap